# Optimal Pure Differentially Private Sparse Histograms in Near-Linear Deterministic Time

### v1

@@ D.1 Protocol
Appendix E Proofs for Section 5
Proof of Theorem 5.4.
$=\sum_{i\geq m}\mu(x_{i})+\sum_{i=1}^{m-1}{\left({\mu(x_{i})-p_{i}}\right)}$
It follows that
$\mathds{TV}({\{{p_{i}}\}},\,{\mu})$ $\leq\frac{1}{2}\cdot{\Bigg{(}{\mu{\left({\overline{\operatorname{C}_{\delta/2}}}\right)}+\lvert{p_{0}-\mu(x_{0})}\rvert+\sum_{i\in[m-1]}\lvert{p_{i}-\mu(x_{i})}\rvert}\Bigg{)}}$ (31)
$\leq\frac{1}{2}\cdot{\Bigg{(}{\mu{\left({\overline{\operatorname{C}_{\delta/2}}}\right)}+\lvert{p_{0}-\mu(x_{0})}\rvert+{({m-1})}\cdot 2^{-\ell}}\Bigg{)}}$ (32)
$\leq\mu{\left({\overline{\operatorname{C}_{\delta/2}}}\right)}+{({m-1})}\cdot 2^{-\ell}$ (33)
$\leq\delta/2+\delta/2$ (34)
$\leq\delta.$ (35)
Space Usage. Clearly, the array $\vec{a}$ requires $\bar{m}\cdot\log_{2}m$ bits of space, where $\bar{m}=2^{{\lceil{\log_{2}m}\rceil}}$. It remains to analyze the space usage of the array $\vec{b}$.
Sampling Bit. Sampling an $I$ uniformly at random from $\bar{m}$ requires exactly $\log_{2}\bar{m}$ bits. Further, since $\vec{b}[I]$ has a binary representation of $\ell-\log_{2}\bar{m}$ bit after the fractional point, sampling $\operatorname{\mathbb{B}ernoulli}\left({\vec{b}[I]}\right)$ can be achieved using less $2$ unbiased random bit in expectation and $\ell-\log_{2}\bar{m}$ in the worst case.
Proof of 5.7.
$\Pr\left[Y=t\right]$ $=\Pr\left[B=0\right]\cdot\Pr\left[S=1\right]\cdot\Pr\left[X=t-1\right]$ (36)
$={\left({1-\Pr\left[\operatorname{\mathbb{DL}ap}\left({p}\right)=0\right]}\right)}\cdot\frac{1}{2}\cdot{\left({1-p}\right)}\cdot p^{t-1}$ (37)
$={\left({1-\frac{1-p}{1+p}}\right)}\cdot\frac{1}{2}\cdot{\left({1-p}\right)}\cdot p^{t-1}$ (38)
$=\frac{1-p}{1+p}\cdot p^{t}.$ (39)
Proof of Lemma 5.8.
For each $t\in\mathbb{N}$,
$\Pr\left[r\cdot X+Y=t\right]$ $=\Pr\left[X={\lfloor{t/r}\rfloor}\right]\cdot\Pr\left[Y=t\mod r\right]$ (40)
$={\left({1-e^{-\varepsilon\cdot r}}\right)}\cdot e^{-\varepsilon\cdot r\cdot{\lfloor{t/r}\rfloor}}\cdot\frac{1-e^{-\varepsilon}}{1-e^{-\varepsilon\cdot r}}\cdot e^{-\varepsilon\cdot{\left({t\mod r}\right)}}$ (41)
$={\left({1-e^{-\varepsilon}}\right)}\cdot e^{-\varepsilon\cdot{\left({r\cdot{\lfloor{t/r}\rfloor}+{\left({t\mod r}\right)}}\right)}}$ (42)
$={\left({1-e^{-\varepsilon}}\right)}\cdot e^{-\varepsilon\cdot t}$ (43)

### v2

@@ Appendix D Circuit-Based MPC Protocol for Sparse Histograms
Appendix E Extension to Add/Remove Model
In Section˜4, we present a deterministic linear-time algorithm achieving optimal $\ell_{\infty}$ error for the pure DP sparse histogram problem under the replacement neighboring model (see Section˜2 for its definition). We now investigate the following question:
Research Question. Can a comparable algorithm be designed under another commonly studied notion of neighboring datasets, the add/remove model, defined formally below?
Add/Remove Neighboring Model.
Two datasets $\mathcal{X}$ and $\mathcal{X}^{\prime}$ are said to be neighboring, denoted $\mathcal{X}\sim\mathcal{X}^{\prime}$, if they differ by the addition or removal of a single entry; that is, $\mathcal{X}^{\prime}=\mathcal{X}\cup{x^{({n+1})}}$ or $\mathcal{X}^{\prime}=\mathcal{X}\setminus{x^{({i})}}$ for some $i\in[n]$.
Accordingly, two histograms $\vec{h},\vec{h}^{\prime}\in\mathbb{N}^{d}$ are called neighboring if they are induced by neighboring datasets under this model.
In the add/remove model, there is no fixed upper bound on the dataset size, so it is often referred to as the unbounded setting. A commonly studied relaxation is the upper-bounded setting, where a public upper bound on the dataset size is assumed.
This add/remove model is fundamentally more challenging than the replacement model—for example, the dataset size $n$ cannot be revealed under add/remove neighbors, whereas it can be published in the replacement model.
Impossibility Result.
The answer to the research question is negative. If an algorithm’s running time depends deterministically on $n$, then its running time leaks $n$: neighboring datasets can have different sizes and thus different execution times. Conversely, if an algorithm’s running time is independent of $n$, then achieving optimal $\ell_{\infty}$ error is impossible. Since $n$ may be arbitrarily larger than the fixed running time, there exist inputs for which the algorithm does not even have enough time to read the entire dataset, making accurate histogram estimation unattainable.
On the other hand, we show that it is possible to design an expected linear-time algorithm with optimal $\ell_{\infty}$ error whose running time remains private, in the sense formalized below.
Time Obliviousness.
The privacy of the running time is captured with by an extended notion differential privacy. Formally, we require the joint distribution of the algorithm’s output and running time to satisfy differential privacy.
Definition E.1 ($({\varepsilon,\delta})$}-DP Time-Oblivious Algorithm (Dov et al., 2023)).
From Replacement to Add/Remove Neighboring Model
Our solution builds on the recent framework of Ratliff and Vadhan (2025) for converting differentially private, time-oblivious algorithms (referred to as Joint Output/Timing private algorithms in their work) from the upper-bounded setting (see the discussion following Definition˜2.2) to the unbounded setting, under the add/remove neighboring model. The original framework is developed in the RAM model, where each memory cell can store an arbitrarily large natural number.
Proposition E.0 ((Ratliff and Vadhan, 2025)).
(46) $\mathds{TV}({\mathcal{M}{({y})}},\,{\mathcal{M}^{\prime}{({y})}})\leq\beta_{1}.$
For clarity of presentation, we tailor the framework to the private sparse histogram problem in Word-RAM model. The pseudocode is given in Algorithm˜8, and the construction relies on two key components.
Algorithm 8 Applying Ratliff and Vadhan’s Framework
2: parameter $\tau\in\mathbb{N}_{+}$ $\triangleright$ as defined in Equation˜2
3: histogram $\mathcal{X}=\{{x^{({1})},\ldots,x^{({n})}}\}$
5:for $k\in\mathbb{N}_{+}$ do
8:  $t\leftarrow\min\{n,\hat{n}_{k}\}$
9:  $\tilde{n}\leftarrow\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({\hat{n}_{k},\varepsilon_{k},\beta_{k}})}}\left({t}\right)$ $\triangleright$ $\tilde{n}\approx\text{clamp}{[{{t+\operatorname{\mathbb{DL}ap}(e^{-\varepsilon_{k}})},{0},{\hat{n}_{k}}}]}$
10:  if $\tilde{n}<\frac{1}{2}\cdot\hat{n}_{k}$ then
11:   Truncate $\mathcal{X}$ to its first $\hat{n}_{k}$ entries if $n>\hat{n}_{k}$
12:   Construct $\vec{h}$ based on $\mathcal{X}$
13:   return $\mathcal{M}_{\textsc{\scriptsize Hist}}(d,\hat{n}_{k},k,a_{\varepsilon},b_{\varepsilon},a_{\gamma},b_{\gamma},\vec{h})$ $\triangleright$ Algorithm˜2
Construction.
First, suppose we have an estimate $\hat{n}$ such that $n\leq\hat{n}\leq c\cdot n$ for some constant $c$. We run Algorithm˜2 with $n$ replaced by $\hat{n}$, keeping all other parameters fixed. This modifies the algorithm in two ways: (1) noise is now sampled over the range $\mathbb{N}_{\leq{\hat{n}}}$, which preserves the original privacy guarantees and tail bounds; and (2) more items are sampled in line 8, which intuitively enhances the privacy of the selected item set $I$. This intuition is formally supported by revisiting the proof of Theorem˜4.1, presented in Section˜4.2. Although replacing $n$ with $\hat{n}$ increases the running time, the asymptotic complexity remains the same, since $\hat{n}\in O(n)$.
Second, Ratliff and Vadhan (2025) show that, for given $\beta_{1}\in(0,1)$, Algorithm˜8 is an $\varepsilon_{1}$-DP time-oblivious algorithm that computes an estimate $\hat{n}$ satisfying $n\leq\hat{n}\leq c\cdot n$ with probability at least $1-\beta_{1}$. The algorithm constructs a sequence of candidate values $\hat{n}_{k}$, starting from a small initial guess and approximately doubling at each step, until it finds some $\hat{n}_{k}\geq n$. At each iteration, it spends privacy budget $\varepsilon_{1}/2^{k}$ to compare $\hat{n}_{k}$ to $n$: it proceeds with probability $1-\beta_{1}/2^{k}$ if $n\geq\hat{n}_{k}$, and halts with the same probability if $n\leq\hat{n}_{k}/2$. The privacy budget and failure probability are geometrically decayed to ensure that their total sums remain bounded by the overall privacy and failure parameters.
There is one caveat: the iteration may terminate with some $\hat{n}_{k}<n$, though this occurs with low probability. If we then run Algorithm˜2 on a histogram constructed from the full dataset $\mathcal{X}$, using $\hat{n}_{k}$ in place of $n=\lvert{\mathcal{X}}\rvert$, the algorithm may fail. In particular, the privacy blanket sampling step (Algorithm˜2, line 8) requires $\hat{n}_{k}+k\geq\lvert{I_{1}}\rvert$, where $\lvert{I_{1}}\rvert$ can be as large as $n$. To address this, Algorithm˜8 truncates $\mathcal{X}$ to its first $\hat{n}_{k}$ elements before constructing $\vec{h}$ and invoking Algorithm˜2. This truncation preserves the sensitivity of neighboring datasets and thus maintains privacy, at the cost of a small utility loss (with low probability).
Privacy, Utility Guarantees, and Running Time.
The joint output/time privacy guarantees and the utility bounds follow directly from Proposition˜A.11 and the properties of Algorithm˜2, as formalized in Theorem˜4.1.
Analyzing the running time of Algorithm˜8 in the word-RAM model requires more care. When $k\in O(\log n)$, the input parameters and internal variables such as $\varepsilon_{k}$, $\beta_{k}$, and $\hat{n}_{k}$ all fit within a constant number of words. If $\mathcal{M}_{\textsc{\scriptsize Hist}}$ (Algorithm˜2) is not invoked, each iteration runs in $\tilde{O}(1/\varepsilon)$ deterministic time, and there are at most $O(\log n)$ such iterations.
When the algorithm terminates, it satisfies $\Pr[n\leq\hat{n}_{k}\leq c\cdot n]\geq 1-\beta_{1}$ for some constant $c$. Thus, with probability at least $1-\beta_{1}$, the cost of invoking $\mathcal{M}_{\textsc{\scriptsize Hist}}$ is $O(n)$.
For larger $k\in\mathbb{N}$ such that $\hat{n}_{k}>c\cdot n$, the variables $\varepsilon_{k}$, $\beta_{k}$, and $\hat{n}_{k}$ may exceed word size and require $O(k)$ bits to represent. This increases the cost of both $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({\hat{n}_{k},\varepsilon_{k},\beta_{k}})}}\left({t}\right)$ and $\mathcal{M}_{\textsc{\scriptsize Hist}}$ by a factor of $O(k)$. However, since the probability of reaching iteration $k$ decays faster than geometrically once $\hat{n}_{k}>c\cdot n$, the overall expected running time remains bounded by $O(n)$.
Appendix F Related Work
F.1. Side-Channel Attacks
Floating Point Attack
Timing Attack
Time-Oblivious Sampling.
As a time-oblivious sampling algorithm can require significantly more random bits than the classical Knuth-Yao sampler (Knuth and Yao, 1976), Dov et al. (2023) introduce a relaxed notion of $\left({\varepsilon,\delta}\right)$} time-oblivious sampler (see Definition˜B.1). They design both $\varepsilon$-pure and $\left({0,\delta}\right)$}-approximate time-oblivious samplers that use logarithmic number of random bits but exponential space.
DP Time-Oblivious Algorithm.
Extending their notion of $(\varepsilon,\delta)$ time-oblivious sampler, Dov et al. (2023) defines DP time-oblivious mechanisms (see Definition˜E.1). They show that any pure DP time-oblivious mechanism over infinitely many datasets gives an irrelevant output with some constant probability. Nevertheless, they demonstrate that a pure DP mechanism can be transformed into a time-oblivious with nearly the same privacy and utility guarantees, although the efficiency of this transformation remains an open problem due to its reliance on manipulating countably many probability values.
In follow-up work, Ratliff and Vadhan (2025) present a new framework for converting pure, time-oblivious differentially private algorithms from the upper-bounded setting (see the discussion following Definition˜2.2) to the unbounded setting, under the add/remove neighboring relation and in the RAM model.
F.2. Noise Generation
  1. (1)
  2. (2)
  3. (3)
    The third has depth $\Theta{\left({\ln{({m+\ell})}}\right)}$, gate count $\Theta{\left({m\ell\cdot\ln{\left({m+\ell}\right)}}\right)}$, and input size $\Theta{\left({m\ln{({m+\ell})}}\right)}$.
F.3. Differentially Private Frequency Estimation
Additional Error Measure
To motivate frequency oracle, we first introduce another commonly considered error measure, the per-query accuracy guarantee.
Definition F.1 ($(\alpha,\beta)$-Accurate Estimator).
A histogram $\tilde{h}$ achieves $(\alpha,\beta)$ per-query accuracy if, for each $i\in[d]$, $\tilde{h}{\left[{i}\right]}$ is an $(\alpha,\beta)$-accurate estimator of $\vec{h}{\left[{i}\right]}$. In the sparse histogram setting, these two notions coincide. Balcer and Vadhan (2019) show that, for any $\beta\in(0,1/2]$, an $\varepsilon$-DP histogram $\tilde{h}$ with $\lVert{\tilde{h}}\rVert_{0}\leq n^{\prime}$ cannot achieve $(\alpha,\beta)$ per-query accuracy for $\alpha\in\Omega{\big({\min\{{\frac{1}{\varepsilon}\ln\frac{d}{n^{\prime}\cdot\beta},\,n}\}}\big)},$ which matches the simultaneous accuracy upper bounds obtained by our algorithm in Section˜4.
There exist other private succinct representations of $\vec{h}$ beyond sparse histograms, known as frequency oracles, which can bypass the per-query accuracy lower bounds discussed below.
Frequency Oracle.
Lebeda et al. (2022) give an $\left({\varepsilon,\delta}\right)$}-DP frequency oracle using $O{\left({n\cdot\ln{\left({d+n}\right)}}\right)}$ bits of space. For $\delta=0$, it achieves expected per-query error $O{\left({1/\varepsilon}\right)}$, query time $O{\left({\ln d}\right)}$, and expected simultaneous error $O{\left({(1/\varepsilon)\cdot\ln d}\right)}$. For $\delta>0$, the per-query error remains $O{\left({1/\varepsilon}\right)}$, the query time becomes $O{\left({\ln{\left({1/\delta}\right)}}\right)}$, and the simultaneous error is $O{\left({(1/\varepsilon)\cdot\ln{\left({1/\delta}\right)}}\right)}$.
As discussed, compared to sparse frequency histograms, these representations typically incur higher computational cost when estimating $\vec{h}{\left[{i}\right]}$ for a given $i\in[d]$. Moreover, they are insufficient for solving the following commonly studied variant of the heavy hitters problem (Bun et al., 2019b; Wu and Wirth, 2022): privately identifying all $i\in[d]$ such that $\vec{h}{\left[{i}\right]}\geq\Delta$, for $\Delta$ as small as possible. Addressing this task with a frequency oracle directly requires querying all $d$ domain elements, incurring $O(d)$ query overhead. Thus, this line of work is orthogonal to our focus on sparse histograms.
Private Histogram.
Cormode et al. (2012) present an $\varepsilon$-DP variant of the stability-based histogram that adds $\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon/2}}\right)$ to all entries of $\vec{h}$ and then releases only those whose noisy counts exceed a threshold $\tau\in\Theta{\left({\frac{1}{\varepsilon}\cdot\ln d}\right)}$. This variant achieves an expected error of $O{\left({\frac{1}{\varepsilon}}\right)}$ for each entry whose true count is at least $\Omega{\left({\frac{1}{\varepsilon}\cdot\ln d}\right)}$, and the expected maximum error over all entries is $O{\left({\frac{1}{\varepsilon}\cdot\ln d}\right)}$. However, a naive implementation of this method still requires $O(d)$ running time. The key observation is that, for entries whose true count is zero, the probability that their noisy count exceeds the threshold $\tau$ is $p_{\tau}\in\tilde{O}{\left({\frac{1}{d}}\right)}$. Therefore, the number of such false positive entries follows the binomial distribution $\operatorname{\mathbb{B}in}({d-\lVert{\vec{h}}\rVert_{0}},{p_{\tau}})$, which has an expected value of $\tilde{O}(1)$, where $\lVert{\vec{h}}\rVert_{0}$ is the number of nonzero entries in $\vec{h}$. Recently, Qiu and Yi (2025) presented a variant of (Cormode et al., 2012) and explicitly discussed how to sample $\operatorname{\mathbb{B}in}({d-\lVert{\vec{h}}\rVert_{0}},{p_{\tau}})$ in the Real-RAM model in $O(1)$ expected time. However, it remains open whether one can sample from $\operatorname{\mathbb{B}in}({d-\lVert{\vec{h}}\rVert_{0}},{p_{\tau}})$ in strict $\tilde{O}(n)$ time in the Word-RAM model.